The test statistic of z=1.36 is obtained when testing the claim that p>0.7

a. Identify the hypothesis test as being​ two-tailed, left-tailed, or​ right-tailed.

b. Find the​ P-value.

c. Using a significance level of alpha = 0.05​, should we reject Upper H 0 or should we fail to reject Upper H 0​?
Answer #1

Solution :

This is the right tailed test .

The null and alternative hypothesis is

H0 : p =0.7

Ha : p > 0.7 upper

= x / n =

P0 =

1 - P0 =

Test statistic = z =1.36

P(z > 1.36) = 1 - P(z <1.36 ) = 1-0.9131=0.0869

P-value = 0.0869

= 0.05

P-value >

fail to Reject the null hypothesis .